# Measuring entanglement without local addressing in quantum many-body simulators via spiral quantum state tomography

### v2

@@ Discussion
We have proposed and validated an efficient method for quantum state reconstruction using compressed sensing based on spiral measurements. This method does not require individual single-qubit addressing, yet our numerical simulations on random target states showed accuracy comparable to that of traditional Pauli measurements. A key advantage of our spiral QST approach is its scalability: different measurement setups with varying spiral pitch angles can be achieved simply by adjusting the application time or magnetic field strength using a single, globally acting magnetic field generator. Unlike conventional Pauli measurements, which require increasingly complex one-by-one qubit addressing as the system size grows, spiral QST remains efficient and practical as quantum platforms continue to scale. This scalability advantage becomes particularly significant as quantum systems expand in size and complexity. Our method is particularly effective for systems where single-qubit addressing is challenging and prohibitively expensive, making it a versatile tool for a broad range of quantum systems.
@@
In addition, the spiral QST method has proven robust against realistic noise sources. When applied to the ground state of specific Hamiltonians, such as the Heisenberg model, our method maintains its effectiveness even in the presence of characteristic experimental inaccuracies. For instance, we have considered scenarios with realistic magnetic field fluctuations and a finite number of measurement repetitions, demonstrating that spiral QST can still faithfully reconstruct the quantum state. This robustness is crucial for practical implementations, where noise and errors are inevitable. Furthermore, the spiral QST approach is suitable for extracting entanglement measures, a key objective of QST in quantum simulations of various fields of physics. The accurate determination of entanglement measures such as von Neumann and Rényi entropies is critical, as these quantities provide deep insights into the quantum correlations and properties of the system.
@@
Despite significant advances, quantum processors are still in a developmental phase, with continuous improvements in error correction and fault tolerance as they progress toward fully scalable and general-purpose quantum computing RevModPhys.95.045005 . During this period of transition, quantum processors are proving their remarkable ability to simulate complex physical systems, demonstrating their practical utility and potential Houck2012-hc ; Daley2022-ok ; Kim2023-nz ; Frey2022-mp ; Mi2022-pl ; Google_Quantum_AI_and_Collaborators2023-bc ; Weimer2010-fu ; Henriet2020-yk ; Gross2017-rb ; Barreiro2011-cp ; Blatt2012-fa ; Monroe2021-tn . The present work not only illustrates the practical application of spiral QST, but also provides a framework for overcoming the experimental difficulties associated with direct measurements of entanglement in large-scale systems. This capability will contribute to a comprehensive understanding of quantum systems and foster interdisciplinary research by bridging condensed matter physics, quantum field theory, gravitational theory, quantum thermodynamics, and other fields through the lens of quantum information science.
Methods
Singular value thresholding
@@ Methods
The quantum compressed sensing protocol gross-11 is aimed at reconstructing a target low-rank density matrix $\rho$ from an under-complete set of tomographic data, namely a set of expectation values of linearly independent operators (measurement operators). Mathematically, the problem is formulated in terms of a matrix variable $\sigma$ as the minimization of the trace norm $||\sigma||_{1}$, subject to ${\rm Tr}(\sigma)=1$ and ${\rm Tr}(\sigma w_{a})={\rm Tr}(\rho w_{a})$, $a=1,\ldots,m$, where $\{w_{a}\}$ are the measurement operators used. We denote the outcome of this problem as $\rho^{\mathrm{rec}}$, the reconstructed density matrix.
In practical numerical calculations, one employs the singular value thresholding (SVT) algorithm Cai2008-zn to minimize $\tau||\sigma||_{1}+||\sigma||^{2}_{2}/2$, where $||A||_{2}\equiv\sqrt{{\rm Tr}(A^{\dagger}A})$ is the Frobenius norm, subject to the above constraints. Clearly this reduces exactly to the minimization of the trace norm for $\tau\to\infty$, but in actual implementations it is enough to take a sufficiently large, but finite, $\tau$; throughout this work $\tau=5$.
@@
Target quantum states are generated as follows. Pure states of dimension $d$ are simply obtained by applying a random unitary matrix drawn from the Haar measure on SU$(d)$ to a fixed reference state. Mixed states of rank $r$ are obtained by first generating a Haar-random pure state in $d\times r$ dimensions and then partial-tracing over the auxiliary $r$ degrees of freedom; this induces a natural measure on the space of rank-$r$ states zyczkowski-01 ; bengtsson-06 ; miszczak-12 . An efficient way to reach the same result is generating a random $d\times r$ matrix $G$ from the Ginibre ensemble  [^1] The Ginibre ensemble is constituted by random complex matrices whose entries (both real and imaginary parts) are i.i.d. variables distributed according to the standard normal distribution $\mathcal{N}(0,1)$. and then defining the target state as $\rho=GG^{\dagger}/{\rm Tr}(GG^{\dagger})$. Finally, the ground states of the Hamiltonians described in the text are calculated by exact diagonalization; the reduced density matrices are then computed via partial tracing over the complement of the subsystem of interest. We have included a component of depolarizing noise to the random states as described in Results to mimic the inaccuracies that may normally occur in the preparation of an actual quantum state; we have not done so for the Hamiltonian ground states and related reduced density matrices, since we intend to focus on the fluctuations of the measurement processes involving application of magnetic field gradients and QGM.
@@
Our SVT codes are tailored to the use of the spiral operators $\{\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\}$ as measurement operators. Let us mention that this set is not orthogonal, but “almost” orthogonal, in the sense that the number of non-orthogonal pairs scales like $1/2^{N}$; it is however linearly independent, except for a few accidental duplications (up to a sign), e.g $\tilde{M}^{XY}(0)=X\otimes\cdots\otimes X=\pm\tilde{M}^{ZX}(\pi)$. These duplications, when they occur, are taken care of before the actual SVT calculation. In each simulation, a certain number $n_{q}$ of pitch angles is chosen from the set in Eq. (4). This choice is random for the analysis of random states, whereas for the Hamiltonian ground states and related reduced density matrices we adopt the strategy of relevant pitch angles described in the following subsection. For each chosen $q$ we use all the non-trivial $12^{N}-$ 1 spiral operators in all of the 3 spin planes. Note that our sets of measurements are therefore very structured, as opposed to the completely random sets pertaining to the original quantum compressed sensing algorithm gross-11 .
@@
The expectation values used in the QST of random quantum states are the subset of $\{{\rm Tr}(\rho\tilde{M}^{\alpha\beta}_{\bm{a}}(q))\}$ dictated by the choice of $q$’s, with the addition of Gaussian fluctuations of width $\sigma=0.1/d$, as described in Results. In the simulation of the optical-lattice setting with QGM, the derivation of the input expectation values deserves a few more words. Upon the repeated preparation of a quantum state, a series of QGM snapshots naturally leads to the measurement of correlations in the $Z$ basis, that is expectation values of the operators $\tilde{M}^{ZX}_{\bm{a}}(0)$. Let us consider that, given any family of operators $\tilde{M}^{\alpha\beta}_{\bm{a}}(q)$, there is a well-defined unitary transformation $U=U_{1}\otimes\cdots\otimes U_{N}$ such that $\tilde{M}^{\alpha\beta}_{\bm{a}}(q)=U\tilde{M}^{ZX}_{\bm{a}}(0)U^{\dagger}$ ($U$ depends on $\alpha\beta$ and $q$, but we omit these indices for simplicity of notation). It follows that
@@
[…]
@@
Random pure (Fig. 2a) 1.5 100 25
Random rank=3 (Fig. 2b) 4 for $n_{q}\geq 12$ 100 25
@@
For the details of the SVT routine itself we refer to the original paper Cai2008-zn . Let us just recall that it is an iterative procedure and its convergence, at least in our applications, is extremely sensitive to the choice of an internal parameter, denoted $\delta$, which can be seen as the “size” of an iteration step. The fine-tuning of $\delta$ as a function of system size, target states and $n_{q}$ is crucial in order to obtain a good $\rho^{\mathrm{rec}}$; our best choices of $\delta$ for each case study are reported in Table 1. The iteration can terminate either if a control condition is met or a maximum number of steps $k_{max}$ is reached (see Table 1). The control condition is $(\sum_{a=1}^{m}({\rm Tr}((\rho^{\mathrm{rec}}-\rho)w_{a}))^{2})^{1/2}<\epsilon$; we take $\epsilon=10^{-1}$. The output $\rho^{\mathrm{rec}}$ in general does not have unit trace (it is typically smaller) and therefore we normalize it a posteriori to have a bona fide density matrix, of which it is possible to calculate fidelity and trace distance with the target state $\rho$. Upon repeating the whole process a number $s$ of times (see Table 1), one can extract a data point and its error bar in Figs. 2-6, corresponding to the mean value and standard deviations of the sampled fidelity and trace distance. As for Figs. 7-8, a sample of $\rho^{\mathrm{rec}}$’s is produced in a similar way and their entanglement entropies are statistically analyzed.
Strategy of relevant pitch angles
@@
For the Heisenberg chain, the spin symmetry of the system leads to additional periodicity and reflection symmetries as a function of $q$. While this behavior is specific to this system, many important quantum states in typical quantum many-body systems exhibit similar symmetries. Based on this observation, it is reasonable to begin QST with $l=0$, then expand sequentially to $l=0,1$, $l=0,1,2$, and so on for such symmetric systems.
On the other hand, in the presence of DM interactions, the tomographically dominant pitch angle shifts to a finite $q$. In this case (see Fig. 9’s caption for the value of the couplings), the contribution of $q=\pi/2$ (which coincides with that of $q=-\pi/2$ due to a specific symmetry) appears to be dominant. Consequently, starting spiral QST from pitch angles around $q=\pi/2$ represents a clever and efficient strategy in such cases.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
acknowledgments
Author contributions
Competing interests
The authors declare no competing interests.

### v3

@@ IV Conclusions
We have proposed and validated an efficient method for quantum state reconstruction using compressed sensing based on spiral measurements. This method does not require individual single-qubit addressing, yet our numerical simulations on random target states showed accuracy comparable to that of traditional Pauli measurements. A key advantage of our spiral QST approach is its scalability: different measurement setups with varying spiral pitch angles can be achieved simply by adjusting the application time or magnetic field strength using a single, globally acting magnetic field generator. Unlike conventional Pauli measurements, which require increasingly complex one-by-one qubit addressing as the system size grows, spiral QST remains efficient and practical as quantum platforms continue to scale. This scalability advantage becomes particularly significant as quantum systems expand in size and complexity. Our method is particularly effective for systems where single-qubit addressing is challenging and prohibitively expensive, making it a versatile tool for a broad range of quantum systems. Also, while the localized spin limit is assumed so far, it is conceivable that the method might be extendable to states with motional degrees of freedom, such as holes and doublons in Hubbard and $t$-$J$ models, which are subjects of intense study in optical-lattice quantum simulators [66, 67, 68, 69].
@@
In addition, the spiral QST method has proven robust against realistic noise sources. When applied to the ground state of specific Hamiltonians, such as the Heisenberg model, it maintains its effectiveness even in the presence of characteristic experimental inaccuracies. For instance, we have considered scenarios with realistic magnetic field fluctuations and a finite number of measurement repetitions, demonstrating that spiral QST can still faithfully reconstruct the quantum state. This robustness is crucial for practical implementations, where noise and errors are inevitable. Furthermore, the spiral QST approach is suitable for extracting entanglement measures, a key objective of QST in quantum simulations of various fields of physics. The accurate determination of entanglement measures such as von Neumann and Rényi entropies is critical, as these quantities provide deep insights into the quantum correlations and properties of the system.
Concerning concrete future directions, it would be interesting to look at the dynamics of entanglement and correlations in various physical situations (quantum quenches, local excitations of quasiparticles, etc.), in particular focusing on the relation between the intrinsic timescales of the target system and those of the measurement process. For example, one can imagine devising an experimental protocol to observe the quantum Mpemba effect in an optical-lattice setup, via the estimation of entanglement asymmetry obtained from the spiral QST of the reduced density matrices [70]. Also, in view of the recent interest for measurement-induced quantum phase transitions, originally formulated within gate-based quantum circuits (for which spiral QST might be a useful tool depending on the platform), one may think of conceptually related phenomena in optical lattices, where the unitary evolution is given by a Hamiltonian and the non-unitary projection by particle losses by inelastic collisions [71], electron beams [72], photoassociation [73], etc. Spiral QST is suited to approach the problem from the area-law side, and could possibly access the shallow volume-law region when the system size is relatively small; it should be interesting to see the dependence of the entanglement entropy (or entanglement negativity for mixed states) on the system parameters, e.g. the dissipation rates.
@@
Despite significant advances, quantum processors are still in a developmental phase, with continuous improvements in error correction and fault tolerance as they progress toward fully scalable and general-purpose quantum computing [74]. During this period of transition, quantum processors are proving their remarkable ability to simulate complex physical systems, demonstrating their practical utility and potential [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. The present work not only illustrates the practical application of spiral QST, but also provides a framework for overcoming the experimental difficulties associated with direct measurements of entanglement in large-scale systems. This capability will contribute to a comprehensive understanding of quantum systems and foster interdisciplinary research by bridging condensed matter physics, quantum field theory, gravitational theory, quantum thermodynamics, and other fields through the lens of quantum information science.
Acknowledgements.
V Author contributions
Appendix A Singular value thresholding
@@
The quantum compressed sensing protocol [21] is aimed at reconstructing a target low-rank density matrix $\rho$ from an under-complete set of tomographic data, namely a set of expectation values of linearly independent operators (measurement operators). Mathematically, the problem is formulated in terms of a matrix variable $\sigma$ as the minimization of the trace norm $\sigma||_{1}$, subject to ${\rm Tr}(\sigma)=1$ and ${\rm Tr}(\sigma w_{a})={\rm Tr}(\rho w_{a})$, $a=1,\ldots,m$, where $\{w_{a}\}$ are the measurement operators used. We denote the outcome of this problem as $\rho^{\mathrm{rec}}$, the reconstructed density matrix.
[…]
@@
Target quantum states are generated as follows. Pure states of dimension $d$ are simply obtained by applying a random unitary matrix drawn from the Haar measure on SU(d)(d) to a fixed reference state. Mixed states of rank $r$ are obtained by first generating a Haar-random pure state in $d\times r$ dimensions and then partial-tracing over the auxiliary $r$ degrees of freedom; this induces a natural measure on the space of rank-$r$ states [87, 88, 89]. An efficient way to reach the same result is generating a random $d\times r$ matrix $G$ from the Ginibre ensemble  [^2] The Ginibre ensemble is constituted by random complex matrices whose entries (both real and imaginary parts) are i.i.d. variables distributed according to the standard normal distribution $\mathcal{N}(0,1)$. and then defining the target state as $\rho=GG^{\dagger}/{\rm Tr}(GG^{\dagger})$. Finally, the ground states of the Hamiltonians described in the text are calculated by exact diagonalization; the reduced density matrices are then computed via partial tracing over the complement of the subsystem of interest. We have included a component of depolarizing noise to the random states as described in Sec. II to mimic the inaccuracies that may normally occur in the preparation of an actual quantum state; we have not done so for the Hamiltonian ground states and related reduced density matrices, since we intend to focus on the fluctuations of the measurement processes involving application of magnetic field gradients and QGM.
@@
Our SVT codes are tailored to the use of the spiral operators $\{\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\}$ as measurement operators. Let us mention that this set is not orthogonal, but “almost” orthogonal, in the sense that the number of non-orthogonal pairs scales like $1/2^{N}$; it is however linearly independent, except for a few accidental duplications (up to a sign), e.g $\tilde{M}^{XY}(0)=X\otimes\cdots\otimes X=\pm\tilde{M}^{ZX}(\pi)$. These duplications, when they occur, are taken care of before the actual SVT calculation. In each simulation, a certain number $n_{q}$ of pitch angles is chosen from the set in Eq. (4). This choice is random for the analysis of random states, whereas for the Hamiltonian ground states and related reduced density matrices we adopt the strategy of relevant pitch angles described in the following subsection. For each chosen $q$ we use all the non-trivial $2^{N}-1$ spiral operators in all of the 3 spin planes. Note that our sets of measurements are therefore very structured, as opposed to the completely random sets pertaining to the original quantum compressed sensing algorithm [21].
@@
The expectation values used in the QST of random quantum states are the subset of $\{{\rm Tr}(\rho\tilde{M}^{\alpha\beta}_{\bm{a}}(q))\}$ dictated by the choice of $q$’s, with the addition of Gaussian fluctuations of width $\sigma=0.1/d$, as described in Sec. II. In the simulation of the optical-lattice setting with QGM, the derivation of the input expectation values deserves a few more words. Upon the repeated preparation of a quantum state, a series of QGM snapshots naturally leads to the measurement of correlations in the $Z$ basis, that is expectation values of the operators $\tilde{M}^{ZX}_{\bm{a}}(0)$. Let us consider that, given any family of operators $\tilde{M}^{\alpha\beta}_{\bm{a}}(q)$, there is a well-defined unitary transformation $U=U_{1}\otimes\cdots\otimes U_{N}$ such that $\tilde{M}^{\alpha\beta}_{\bm{a}}(q)=U\tilde{M}^{ZX}_{\bm{a}}(0)U^{\dagger}$ ($U$ depends on $\alpha\beta$ and $q$, but we omit these indices for simplicity of notation). It follows that
@@
This means that the new expectation values can be obtained by operating on the target state with $U^{\dagger}$, namely $U^{\dagger}\rho U\equiv\rho^{\prime}$. This is exactly what the optical (Rabi pulses) and magnetic (field gradient pulses) operations achieve. We include the Gaussian fluctuations adopted for the zero point of the magnetic field (see Sec. III) into all our numerical implementations of $U$. At this stage, the diagonal elements of $\rho^{\prime}$ simply give a probability distribution for the QGM snapshots, e.g. $\rho^{\prime}_{11}$ is the probability of obtaining $|\uparrow\cdots\uparrow\rangle$, $\rho^{\prime}_{22}$ is the probability of obtaining $|\uparrow\cdots\uparrow\downarrow\rangle$, etc. In the numerical simulations, we draw a sufficient number (denoted “reps”) of snapshots from this distribution to evaluate each family of expectation values $\langle\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\rangle$. In statistics and probability theory, “reps” is basically the sample complexity of learning the distribution; for a fixed distance to the true distribution (measured by the $l_{1}$ distance, the classical equivalent of the trace distance), it is expected to scale linearly with $d$ [91]. This is compensated by the fact that the number of obtained independent expectation values also has the same scaling.
@@
Random pure (Fig. 2(a)) 1.5 100 25
Random rank=3 (Fig. 2(b)) 4 for $n_{q}\geq 12$ 100 25
@@
For the details of the SVT routine itself we refer to the original paper [43]. Let us just recall that it is an iterative procedure and its convergence, at least in our applications, is extremely sensitive to the choice of an internal parameter, denoted $\delta$, which can be seen as the “size” of an iteration step. The fine-tuning of $\delta$ as a function of system size, target states and $n_{q}$ is crucial in order to obtain a good $\rho^{\mathrm{rec}}$; our best choices of $\delta$ for each case study are reported in Table 1. A preliminary indicator of the goodness of $\delta$ is the trace of the reconstructed matrix being close to one (before the final normalization) and having a limited spread around the mean value over various simulations. The iteration can terminate either if a control condition is met or a maximum number of steps $k_{max}$ is reached (see Table 1). The control condition is $(\sum_{a=1}^{m}({\rm Tr}((\rho^{\mathrm{rec}}-\rho)w_{a}))^{2})^{1/2}<\epsilon$; we take $\epsilon=10^{-1}$. The output $\rho^{\mathrm{rec}}$ in general does not have unit trace (it is typically smaller) and therefore we normalize it a posteriori to have a bona fide density matrix, of which it is possible to calculate fidelity and trace distance with the target state $\rho$. Upon repeating the whole process a number $s$ of times (see Table 1), one can extract a data point and its error bar in Figs. 2-6, corresponding to the mean value and standard deviations of the sampled fidelity and trace distance. As for Figs. 7-8, a sample of $\rho^{\mathrm{rec}}$’s is produced in a similar way and their entanglement entropies are statistically analyzed.
Appendix B Strategy of relevant pitch angles
@@
For the Heisenberg chain, the spin symmetry of the system leads to additional periodicity and reflection symmetries as a function of $q$. While this behavior is specific to this system, many important quantum states in typical quantum many-body systems exhibit similar symmetries. Based on this observation, it is reasonable to begin QST with $\{l=0\}$, then expand sequentially to $\{l=0,1,-1\}$, $\{l=0,1,-1,2,-2\}$, and so on for such symmetric systems. It is interesting to note that, if we also include the sets $\{0,1\},\{0,1,-1,2\}$, etc., which do not respect the symmetry, we encounter instances of non-monotonicity of fidelity and trace distance as a function of $m/d^{2}$. This indicates, to some extent unexpectedly, that the algorithm is indeed quite sensitive to whether or not the choice of $q$’s is symmetric. Incidentally, the above sequence of symmetric $q$-sets is used also for the reduced density matrix (see Fig. 6), which turns out to have an approximate inversion symmetry.
In the presence of DM interactions, the tomographically dominant pitch angle shifts to a finite $q$. In this case (see Fig. 9’s caption for the value of the couplings), the contribution of $q=\pi/2$ (which coincides with that of $q=-\pi/2$ due to a specific symmetry) appears to be dominant. Consequently, starting spiral QST from pitch angles around $q=\pi/2$ represents a clever and efficient strategy in such cases, but the question is how to add the subsequent pitch angles. One may choose to progressively add the adjacent ones in the following fashion: (for $N=8$) $\{l=4\},\{4,5\},\{4,5,3\},\{4,5,3,6\},\{4,5,3,6,2\}$, etc.. This is reminiscent of the sequence for the Heisenberg case, with $l=4$ in the role of $l=0$. Hover this leads to a tomographic sequence that is worse overall compared to that in Fig. 5 (except for the first data point, as argued above). One possible way to effectively improve the spiral QST for the DM case even further is to modify the algorithm so that the pitch angle of the measurement operators in the $XY$ (where the spin spiral characterizing the ground state develops) is independent of those in the $YZ$ and $ZX$ plane.